1.5x-2y=-0.25
3x+1.5y=6.375
Solve with addition method

Answer by nabla(475)

First, multiply the first equation by -2.
giving
-3x+4y=.5
3x+1.5y=6.375
Now we can add them.
5.5y=6.875
y=1.25

And x must be
1.5x-2(1.25)=-.25
1.5x=2.25
x=1.5
check:
1.5(1.5)-2(1.25)=-.25
3(1.5)+1.5(1.25)=6.375